# Learning finitely correlated states: stability of the spectral reconstruction

### v2

@@ Appendix C Proofs for the non-homogeneous case
$U_{j}^{\intercal}\mathcal{F}_{j}(\mathcal{E}^{[j+1,j+r]}(A\otimes\mathds{1}_{\mathcal{B}}))$ $=U_{j}^{\intercal}\Omega^{[j-l+1,j,j+r]}[A]=D_{j}O_{j}[A]\neq 0.$ (217)
Then we can prove Proposition 5.8
Proof.
We have, for any $X_{j+1}\otimes\cdots\otimes X_{j+r}\in\mathcal{A}_{[j+1,j+r]}$, $X_{j-l-1}\otimes\cdots\otimes X_{j}\in\mathcal{A}_{[j-l+1,j]}$,
$\mathcal{F}_{j}[\mathcal{E}_{X_{j+1}}^{(j+1)}\cdots{\mathcal{E}_{X_{j+r}}^{(j+r)}}(\mathds{1})](X_{j-l+1}\otimes\cdots\otimes X_{j})=\rho_{j-l}\mathcal{E}_{X_{j-l+1}}^{(j-l-1)}\cdots\mathcal{E}_{X_{j}}^{(j)}\mathcal{E}_{X_{(j+1)}}^{(j+1)}\cdots{\mathcal{E}_{X_{j+r}}^{(j+r)}}(\mathds{1}),$ (218)
in particular
$\mathcal{F}_{j}[\mathcal{E}_{X_{j+1}}^{(j+1)}\cdots{\mathcal{E}_{X_{(j+r)}}^{(j+r)}}(\mathds{1})](X_{j-l+1}\otimes\cdots\otimes X_{j})=\Omega^{[j-l+1,j,j+r]}[X_{j+1}\otimes\cdots\otimes X_{j+r}](X_{j-l+1}\otimes\cdots\otimes X_{j}),$ (219)
and
$\mathcal{F}_{j-1}[\mathcal{E}_{X_{j}}^{(j)}\mathcal{E}_{X_{j+1}}^{(j+1)}\cdots{\mathcal{E}_{X_{k}}^{(j+r)}}(\mathds{1})](X_{j-l}\otimes\cdots\otimes X_{j-1})=\Omega^{[j-l,j-1,j+r]}_{X_{j}}[X_{j+1}\otimes\cdots\otimes X_{j+r}](X_{j-l}\otimes\cdots\otimes X_{j-1}),$ (220)
@@
$\mathbb{K}^{(j)}_{A}(x)$ $:=U_{j-1}^{\intercal}\mathcal{F}_{j-1}\mathcal{E}^{[j,j+r]}(A\otimes(U^{\intercal}_{j}\Omega^{[j-l+1,j,j+r]})^{+}(x)\otimes\mathds{1}_{\mathcal{B}})$ (221)
$=U_{j-1}^{\intercal}\mathcal{F}_{j-1}\mathcal{E}^{(j)}_{A}(\mathcal{E}^{[j+1,j+r]}((U^{\intercal}_{j}\Omega^{[j-l+1,j,j+r]})^{+}(x)\otimes\mathds{1}_{\mathcal{B}}))$ (222)
$=U_{j-1}^{\intercal}\mathcal{F}_{j-1}\mathcal{E}^{(j)}_{A}((U^{\intercal}_{j}\mathcal{F}_{j})^{-1}U^{\intercal}_{j}\mathcal{F}_{j}\mathcal{E}^{[j+1,j+r]}((U_{j}^{\intercal}\Omega^{[j-l+1,j,j+r]})^{+}(x)\otimes\mathds{1}))$ (223)
@@
Appendix D Guarantees on marginal size for translation invariant states
We observe that if the rank $m$ of the finitely correlated state viewed as a bilinear form is known, the construction can be made completely explicit. First, we have the following lemma.
Lemma D.1.
Define $V_{i}$ the vector space generated by functionals $\Psi_{X}:\mathcal{A}_{L}\rightarrow\mathbb{C}$ of the form
$\Psi_{X}(A_{1}\otimes\cdots\otimes A_{n}):=\omega(X\otimes A_{1}\otimes\cdots\otimes A_{n}),$ (225)
with $X\in\mathcal{A}_{[-i+1,0]}$ and $n\in\mathbb{N}$. If $\omega$ has a minimal realization of dimension $m$, then there exists $i\leq m$ such that $V_{i}=V_{j}=\Psi(\mathcal{A}_{R})$ for each $j\geq i$. In particular, this holds for $V_{m}$.
@@ Appendix D Guarantees on marginal size for translation invariant states
Let $i$ be such that $\dim V_{i}=\dim V_{i+1}$, i.e. $V_{i}=V_{i+1}$. Let $\Psi_{Y_{k}}$, $k=1,\dots,\dim V_{i}$, be a basis of $V_{i}$, so $Y_{k}\in\mathcal{A}_{[-i+1,0]}$. If $X=\sum_{j}X_{j}\otimes A_{j}\in\mathcal{A}_{[-i-1,-1]}\otimes\mathcal{A}=\mathcal{A}_{[-i-1,0]}$, i.e. $\Psi_{X}\in V_{i+2}$, then for $A\in\mathcal{A}_{R}$ using translation invariance of the FCS $\omega$:
$\Psi_{X}(A)=\sum_{j}\Psi_{X_{j}}(A_{j}\otimes A)=\sum_{j,k}c_{jk}\Psi_{Y_{k}}(A_{j}\otimes A)=\sum_{j,k}c_{jk}\Psi_{Y_{k}\otimes A_{j}}(A)=\Psi_{Y}(A)$
where $Y=\sum_{j,k}c_{jk}Y_{k}\otimes A_{j}\in\mathcal{A}_{[-i,0]}$. Therefore, $V_{i+2}=V_{i+1}$ and by induction $V_{i}=V_{j}=\Psi(\mathcal{A}_{L})$ for all $j\geq i$. Because of $\dim(\Psi(\mathcal{A}_{L}))=m$ by assumption, we have $\dim V_{i}=m$ and thus $i\leq m$ since $\dim V_{j}$ has to increase monotonically in $j$. ∎
Let $\Phi:\mathcal{A}_{L}\to\mathcal{A}_{[1,m]}^{*}$ be the map such that $\Phi_{X}(Y)=\Psi_{Y}(X)$ where $X\in\mathcal{A}_{L},Y\in\mathcal{A}_{[1,m]}$. Furthermore, $\Phi(\mathcal{A}_{L})\cong\Psi(\mathcal{A}_{R})^{*}\cong\Psi(\mathcal{A}_{[1,m]})^{*}$, and vice versa, the last being through because of the above Lemma. The argument in the proof of the above Lemma can also be applied to the analogous spaces $W_{i}$ for $\Phi$ showing that $\Phi(\mathcal{A}_{L})\cong\Phi(\mathcal{A}_{[-m+1,0]})$.
Therefore, it suffices to parametrize $\Psi_{X}$ by $X\in\mathcal{A}_{[1,m]}$ and restrict to $\mathcal{A}_{[-m+1,0]}$ so one has
Corollary D.2.
$\Psi_{X}(A):=\omega(X\otimes A),$ (226)
Appendix E Normalized Gell-Mann matrices
@@
$\lambda_{6}=\frac{1}{\sqrt{3}}\begin{pmatrix}0&0&0\\ 0&0&1\\ 0&1&0\end{pmatrix}\quad\lambda_{7}=\frac{1}{\sqrt{2}}\begin{pmatrix}0&0&0\\ 0&0&-i\\ 0&i&0\end{pmatrix}\quad\lambda_{8}=\frac{1}{\sqrt{6}}\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&-2\end{pmatrix}$

### v3

@@ D.2 Bound for $C^{*}$-$(r,l,m)$-FCS
$U_{j}^{\intercal}\mathcal{F}_{j}(\mathcal{E}^{[j+1,\min(j+r,N)]}(A\otimes\mathds{1}_{\mathcal{B}}))$ $=U_{j}^{\intercal}\Omega^{[j-l+1,j,j+r]}[A]=D_{j}O_{j}[A]\neq 0.$ (227)
Then we can prove Proposition D.8.
Proof of Proposition D.8.
We have, for any $X_{j+1}\otimes\cdots\otimes X_{\min(j+r,N)}\in\mathcal{A}_{[j+1,\min(j+r,N)]}$, $X_{\max(j-l+1,1)}\otimes\cdots\otimes X_{j}\in\mathcal{A}_{[\max(j-l+1,1),j]}$,
$\mathcal{F}_{j}[\mathcal{E}_{X_{j+1}}^{(j+1)}\cdots{\mathcal{E}_{X_{\min(j+r,N)}}^{(\min(j+r,N))}}(\mathds{1})](X_{\max(j-l+1,1)}\otimes\cdots\otimes X_{j})$
$=\rho_{j-l}\mathcal{E}_{X_{\max(j-l+1,1)}}^{\max(j-l+1,1)}\cdots\mathcal{E}_{X_{j}}^{(j)}\mathcal{E}_{X_{(j+1)}}^{(j+1)}\cdots{\mathcal{E}_{X_{\min(j+r,N)}}^{(\min(j+r,N))}}(\mathds{1}_{\mathcal{B}}),$ (228)
in particular, for $X\in\mathcal{A}_{[j+1,\min(j+r,N)]}$, $Y\in\mathcal{A}_{[\max(j-l+1,1),j]}$
$\mathcal{F}_{j}[\mathcal{E}^{[j+1,\min(j+r,N)]}(X\otimes\mathds{1}_{\mathcal{B}})](Y)=\Omega^{[j-l+1,j,j+r]}[X](Y),$ (229)
and for $X\in\mathcal{A}_{[j+1,\min(j+r,N)]}$, $X_{j}\in\mathcal{A}_{[j,j]}$,$Y\in\mathcal{A}_{[\max(j-l,1),j-1]}$
$\mathcal{F}_{j-1}[\mathcal{E}^{[j,\min(j+r,N)]}(X_{j}\otimes X\otimes\mathds{1}_{\mathcal{B}})](Y)=\Omega^{[j-l,j-1,j+r]}_{X_{j}}[X](Y).$ (230)
@@
$\mathbb{K}^{(j)}_{A}(x)$ $:=U_{j-1}^{\intercal}\mathcal{F}_{j-1}\mathcal{E}^{[j,\min(j+r,N)]}(A\otimes(U^{\intercal}_{j}\Omega^{[j-l+1,j,j+r]})^{+}(x)\otimes\mathds{1}_{\mathcal{B}})$ (231)
$=U_{j-1}^{\intercal}\mathcal{F}_{j-1}\mathcal{E}^{(j)}_{A}(\mathcal{E}^{[j+1,\min(j+r,N)]}((U^{\intercal}_{j}\Omega^{[j-l+1,j,j+r]})^{+}(x)\otimes\mathds{1}_{\mathcal{B}}))$ (232)
$=U_{j-1}^{\intercal}\mathcal{F}_{j-1}\mathcal{E}^{(j)}_{A}((U^{\intercal}_{j}\mathcal{F}_{j})^{-1}U^{\intercal}_{j}\mathcal{F}_{j}\mathcal{E}^{[j+1,\min(j+r,N)]}((U_{j}^{\intercal}\Omega^{[j-l+1,j,j+r]})^{+}(x)\otimes\mathds{1}_{\mathcal{B}}))$ (233)
@@
As an immediate consequence, we have that the maps $M_{j-1}^{-1}{\mathbb{K}}^{(j)}_{(\cdot)}M_{j}$ with $M_{j}={U}_{j}^{\intercal}\mathcal{F}_{j}R_{j}$ for $12\leq j\leq N-$ 1, $M_{0}(z)=z$, $M_{N}(z)=z$, are completely positive and unital with respect to the orders on $\mathcal{V}^{\prime}_{j}$ and $\mathcal{V}^{\prime}_{j-1}$, and the error propagation bound from Theorem D.5 applies. Analogously to the translation invariant case, we have that
By using the error progapagation bound in Theorem D.5, we obtain
Theorem D.10.
Let $\omega$ be a $C^{*}$, $(l,r,m)$-finitely correlated state, with associated maps $\Omega^{[i,j,k]}$ satisfying $\eta\leq\min_{j=0,\dots,N}\sigma_{m}(\Omega^{[j-l+1,j,j+r]})/2$. Assume we have estimates of marginals $\omega_{[\max(j-l,1),\min(j+r,N)]}$, $\omega_{[\max(j-l+1,1),\min(j+r,N)]}$, $j=0,...,N$, such that the Hilbert-Schmidt norm errors are smaller than $\frac{\epsilon\eta^{3}}{20td_{\mathcal{B}}\sqrt{d_{\mathcal{A}}}}$. The estimated realization parameters from LearnFCS(l,r) (given by D.4) give rise to an estimated operator $\hat{\omega}$ such that
$\frac{\|\hat{\omega}-\omega\|_{1}}{2}\leq\epsilon,$ (236)
Appendix E Learning states $\epsilon$-close to FCS
A feature of our learning algorithm is that it works even if the true state is sufficiently close to a finitely correlated state. The following proposition holds.
Proposition E.1.
Suppose that a state $\sigma$ is close to an $(l,r,m)$-finitely correlated $\omega$, say $d_{\mathrm{Tr}}(\sigma_{t},\omega_{t})\leq\xi(t)$, for any marginals of size $t$, and that estimates of the marginals in Hilbert-Schmidt distance at precision $\epsilon_{HS}$ are given to the reconstruction algorithm of Theorem D.6. Then, if $\xi(l+r)\leq\sigma_{m_{j}}(\Omega^{[j-l+1,j,j+r]})/8$ for $j=0,\dots,N$, the estimate $\hat{\sigma}$ obtained from the reconstruction algorithm satisfies $\frac{\|\sigma-\hat{\sigma}\|_{1}}{2}\leq\xi(N)+\frac{\|\omega-\hat{\sigma}\|_{1}}{2}$, with $\frac{\|\omega-\hat{\sigma}\|_{1}}{2}<D\epsilon$, where $D>0$ is a universal constant and $\epsilon_{HS}+\xi(l+r+1)=\min_{j=0,\dots,N}\frac{\epsilon\sigma^{3}_{m_{j}}(\Omega^{[j-l+1,j,j+r]})}{20Nm\sqrt{d_{\mathcal{A}}}}$.
@@ Appendix E Learning states $\epsilon$-close to FCS
Observe that $\|\hat{\Omega}^{[i,j,k]}-{\Omega}^{[i,j,k]}\|_{2}\leq\epsilon_{HS}+\xi(k-i+1)$ for the relevant marginals. Provided that $\xi(l+r)\leq\sigma_{m_{j}}(\Omega^{[j-l+1,j,j+r]})/8$, then the truncated singular value decompositions return maps $\hat{U}_{j}$ of rank $m_{j}$ and $\|\hat{\Omega}^{[i,j,k]}-{\Omega}^{[i,j,k]}\|_{2}\leq\sigma_{m_{j}}({\Omega}^{[i,j,k]})/3$ for the relevant marginals. Therefore the error propagation bound can be invoked. ∎
Many physically relevant many-body quantum states can be efficiently approximated by finitely correlated states. One such class corresponds to states that are efficiently approximated by quantum circuits: for an $n=2t$ qudit system initiated in the tensor product state $|0\rangle^{\otimes n}$, consider the following possibly non-unitary depth $D$ quantum circuit with brickwork architecture (see left part of Figure 6 for an illustration in the case of $D=2$):
$\mathcal{C}:=\prod_{\ell=1}^{D}\,\Phi^{(\ell)}\,,$
$\Phi^{(\ell)}:=\bigotimes_{j=1}^{t}\Phi^{(\ell)}_{j}\,,$
$\mathcal{C}:=\prod_{j=1}^{t}\Psi_{j}\,.$
An important class of many-body quantum states which are known to be approximated by outputs of (non-unitary) circuits as described above, namely Gibbs states of one-dimensional local Hamiltonians. More in general, approximation of Gibbs states in terms of matrix product density operators have been studied [8, 9, 10, 36]. The result of Proposition E.1 suggests a path for learning one-dimensional Gibbs states in terms of their approximations by matrix product operators. The recent years have seen the development of many learning algorithms for Gibbs states  [41, 50, 51, 52, 4, 53, 54, 55]. These approximations involve blocking sites of the original chain, so that the local dimension of the new chain is $d_{\mathcal{A}}=d^{L}$ for some $L$ equal to the size of each block of sites and $d$ the original site dimension. To give learning guarantees via these results together with Proposition E.1 we would need to
  • (i) The smallest singular values $\sigma_{m_{j}}({\Omega}^{[j-l+1,j,j+r]})$ of the MPO approximations are large compared with the approximation error. At present we do not have any https://optimization-online.org/wp-content/uploads/2018/01/6441.pdway to prove bounds for these quantities for MPO approximations of Gibbs states or for outputs of quantum circuits. We only have the upper bound $\sigma^{2}_{m_{j}}({\Omega}^{[j-l+1,j,j+r]})\leq\operatorname{Tr}[({\Omega}^{[j-l+1,j,j+r]})^{\dagger}{\Omega}^{[j-l+1,j,j+r]}]/m=\operatorname{Tr}[\omega_{[j-l+1,j+r]}^{2}]/m$.
  • (ii) To make the final trace norm bound in Proposition E.1 non-trivial, the quantity
    $\min_{j=0,\dots,N}\frac{\sigma^{3}_{m_{j}}(\Omega^{[j-l+1,j,j+r]})}{20Nm\sqrt{d_{\mathcal{A}}}}$ must also be sufficiently large compared to the approximation error $\xi(l+r+1)$, otherwise $D\epsilon$ may be larger than $1$.
  • (iii) check that the MPO approximations are $(l,r,m)$-finitely correlated states for sufficiently small $l,r$. This condition is not easy to verify even if the MPO is a circuit. However, it is reasonable to believe that it holds generically as long as $l$ and $r$ are $\Omega(\log m)$; therefore, if it does not hold from the start, it can be verified by an arbitrary small perturbation of the MPO approximation. A formal analysis of this fact and of how the perturbation affects the other conditions may be feasibile. Since the other conditions are anyway hard to verify, we avoid pursuing this.
We mention that the positive MPO approximation in [36] has values of $m$ and $L$ which are compatible with conditions $(i)$ and $(ii)$ being satisfied, but since we do not know how large $l$ and $r$ should be, and we cannot lower bound $\sigma_{m_{j}}(\Omega^{[j-l+1,j,j+r]})$, we do not explore this further.
These observations suggests that to guarantee that spectral reconstruction algorithm work for Gibbs states, we need either to improve our analysis of MPO approximations or to modify the algorithm and its analysis. We leave these questions to future work.
In contrast, it was shown in [56, Theorem 1.3] that learning classical Gibbs measures with $\epsilon$ average recovery in trace distance requires at least $n/\epsilon^{2}$ samples. Since classical Gibbs states are finitely correlated states, the latter directly implies the following:
Proposition E.2.
Appendix F Numerical experiments
F.1 The model: ground state of the AKLT Hamiltonian
that is completely defined by
where |±$|\pm\frac{1}{2}\rangle$ and $|1\rangle,|0\rangle,|-1\rangle$ denote orthonormal bases of $\mathbb{C}^{2}$ and $\mathbb{C}^{3}$ respectively. Then the states $w_{\theta}$ that we consider are given by the realization $(\mathcal{V},e,\mathbb{E}_{\theta},\rho)$ described by,
  • $\mathcal{V}=\mathbb{M}_{2}$.
  • $\rho(B)=\frac{1}{2}\operatorname{Tr}(B)$ for $B\in\mathbb{M}_{2}$.
  • $e=\mathds{1}_{\mathbb{M}_{2}}\in\mathbb{M}_{2}$.
F.2 Results
For the simulations, we restrict our attention to the ground state of the AKLT Hamiltonian (241) and fix $\cos\theta=\sqrt{\frac{2}{3}}$ in the realization of the previous Section F.1. In order to fix a basis for the one site algebra $\mathbb{M}_{3}$ we use the normalized Gell-Mann matrices $\{\lambda_{i}\}_{i=0}^{8}$
@@ F.2 Results
$\lambda_{6}=\frac{1}{\sqrt{3}}\begin{pmatrix}0&0&0\\ 0&0&1\\ 0&1&0\end{pmatrix}\quad\lambda_{7}=\frac{1}{\sqrt{2}}\begin{pmatrix}0&0&0\\ 0&0&-i\\ 0&i&0\end{pmatrix}\quad\lambda_{8}=\frac{1}{\sqrt{6}}\begin{pmatrix}1&0&0\\ 0&1&0\\ 0&0&-2\end{pmatrix},$
where $\lambda_{0}$ is the normalized identity such that $\operatorname{Tr}(\lambda_{i}\lambda_{j})=\delta_{i,j}$. We are interested in recovering the reduced density matrices of $k$ contiguous sites that are described by
$\rho^{(k)}_{\theta}=\sum_{i_{1},\dots,i_{k}}\omega_{\theta}(\lambda_{i_{1}}\otimes\cdots\otimes\lambda_{i_{k}})\lambda_{i_{1}}\otimes\cdots\otimes\lambda_{i_{k}},$ (243)
where $i_{1},\dots,i_{k}\in\{0,1,\dots,8\}^{\otimes k}$. For the observable regular realization (Proposition A.5), we fix the right and left finite-dimensional unital subalgebras $\mathcal{C}_{R}=\mathcal{C}_{L}=\mathbb{M}_{3}$ for two contiguous sites with basis $\{\lambda_{i}\}_{i=0}^{8}$ for both $\mathcal{C}_{R}$ and $\mathcal{C}_{L}$. In order to test our algorithm, we will use some simulated data. Specifically, since we have access to the realization (F.1) we can compute the maps $\Omega$ and $\Omega_{A}$ for $A\in\{\lambda_{i}\}_{i=0}^{8}$ which are the observables that we want to read from the experiment. For completeness, we numerically check that $\text{rank}(\Omega)=4$ which is consistent with the realization given in Section F.1 since $\text{dim}(\mathcal{V})=4$. Then we fix the errors $\epsilon,\epsilon^{\prime}>0$ and generate the simulated observables $\hat{\Omega}$ and $\hat{\Omega}_{A}$ as follows
for all $i=0,\dots,8$ and $P,P_{i}^{\prime}\in\mathbb{R}^{9\times 9}$ are matrices where each entry is drawn from a Gaussian distribution with zero mean and unit variance. With this construction, we can fix the errors $\|\Omega-\hat{\Omega}\|_{2}$ and $\|\Omega_{(\cdot)}-\hat{\Omega}_{(\cdot)}\|_{2}$. Given the singular value decomposition of $\hat{\Omega}$ we need to achieve the invertibility condition of Theorem C.3 that says that $\hat{U}^{\intercal}U$ is invertible. In general $\hat{U}^{\intercal}\in\mathbb{R}^{9\times 9}$ since the perturbation $P$ will make $\hat{\Omega}$ full rank. In order to avoid this issue we can use as input rank $m=4$ and we truncate $\hat{U}$ such that we keep the columns corresponding to the 4 largest singular values of $\hat{\Omega}$. Another possibility is to use Lemma G.1 in order to justify that the singular values with absolute values smaller than a threshold $\eta/2$ can be considered 0s. Thus, setting $\eta$ small enough following conditions in Definition 2.2 we can truncate $\hat{U}$ keeping the columns corresponding to singular values greater than the threshold $\eta/2$. We do not simulate estimates of $\Omega(\mathds{1})$ and $\tau\Omega$, as we use the corresponding values of $\Omega$. This is fine for the purpose of this numerical illustration, as $\Omega$ acts on just two sites and thus the Hilbert-Schmidt error on $\widehat{\Omega(\mathds{1})}$ and $\widehat{\tau\Omega}$ computated from $\hat{\Omega}$ is of the same order of $\|\Omega-\hat{\Omega}\|_{2}$.
Using the simulated observables $\hat{\Omega},\hat{\Omega}_{\lambda_{i}}$ and the truncated $\hat{U}$ we use the spectral state decomposition of Definition C.1 and compute the correlations
Appendix G Matrix perturbation theory
The following perturbation bounds can be found in [57]. Note that the original reference uses the convention of matrices with more rows than columns, while we need the opposite one in our analysis. This is not an issue as the norms and the Moore-Penrose pseudo-inverse are respectively invariant and commuting with transposition.
Lemma G.1 (Theorem 4.11, in [57], p.204).
Let $A,\tilde{A}\in\mathbb{R}^{m\times n}$ with $n\geq m$. Then if $\sigma_{1}\geq\cdots\geq\sigma_{n}$ and $\tilde{\sigma}_{1}\geq\cdots\geq\tilde{\sigma}_{n}$ are the singular values of $A$ and $A^{\prime}$ respectively then
$|\sigma_{i}-\tilde{\sigma}_{i}|\leq\|A-\tilde{A}\|_{2\rightarrow 2}\quad\text{for}\quad i=1,\dots,n.$ (246)
Lemma G.2 (Theorem 3.8 in [57], p.143).
We also need the following results.
Lemma G.3 (Corollary 17 of [27]).
Let $A\in\mathbb{R}^{m\times n}$, with $m\geq n$, have rank $k\leq n$, and let $U\in\mathbb{R}^{m\times k}$ be the matrix of $n$ left singular vectors corresponding to the non-zero singular values $\sigma_{1}\geq\ldots\geq\sigma_{k}>0$ of $A$. Let $\tilde{A}=A+E$. Let $\tilde{U}\in\mathbb{R}^{m\times k}$ be the matrix of $k$ left singular vectors corresponding to the largest $k$ singular values $\tilde{\sigma}_{1}\geq\ldots\geq\tilde{\sigma}_{n}$ of $\tilde{A}$, and let $\tilde{U}_{\perp}\in\mathbb{R}^{m\times(m-k)}$ be the remaining left singular vectors. Assume $\|E\|_{2\rightarrow 2}\leq\epsilon\sigma_{k}$ for some $\epsilon<1$. Then:
  1. 1.
    $\tilde{\sigma}_{k}\geq(1-\epsilon)\sigma_{k}$,
  2. 2.
    $\|\tilde{U}_{\perp}^{\intercal}U\|_{2\rightarrow 2}\leq\|E\|_{2\rightarrow 2}/\tilde{\sigma}_{k}$.
Lemma G.4.
$\sigma_{d_{1}}(S)\sigma_{d_{2}}(T)\leq\frac{\sigma_{i}(SAT)}{\sigma_{i}(A)}\leq\sigma_{1}(S)\sigma_{1}(T)$ (248)
Lemma G.5.
In the notations of Section C, suppose $\|\Omega-\hat{\Omega}\|_{2\rightarrow 2}\leq\varepsilon\cdot\sigma_{m}(\Omega)$ for some $\varepsilon<1/2$. Then $\varepsilon_{0}:=\|\Omega-\hat{\Omega}\|_{2\rightarrow 2}^{2}/((1-\varepsilon)\sigma_{m}(\Omega))^{2}<1$ and:
  1. 1.
  2. 2.
    $\sigma_{m}(\hat{U}^{\intercal}U)\geq\sqrt{1-\varepsilon_{0}}$,
  3. 3.
Proof.
Since $\sigma_{m}(\hat{U}^{\intercal}\hat{\Omega})=\sigma_{m}(\hat{\Omega})$, the (1.) is immediate from Lemma G.3. (2.) follows since for any $x\in\mathbb{C}^{m}$, $\|\hat{U}^{\intercal}Ux\|_{2}\geq\|x\|_{2}\sqrt{1-\|\hat{U}_{\perp}^{\intercal}U\|_{2\rightarrow 2}^{2}}\geq\|x\|_{2}\sqrt{1-\varepsilon_{0}}$ by Lemma G.3 and the fact that $\varepsilon_{0}<1$ (see Lemma 21 of [27] for a full proof). (3.) follows from Lemma G.4 noticing that $\sigma_{m}(\hat{U}^{\intercal}\Omega)=\sigma_{m}((\hat{U}^{\intercal}U)(U^{\intercal}\Omega))$, $\sigma_{m}(U^{\intercal}\Omega)=\sigma_{m}(\Omega)$, and choosing $S=\hat{U}^{\intercal}U$, and $A$ obtained from padding $D$ with zeros and $T$ a unitary which extends $O$. ∎